OpenStudy (anonymous):

Solve (x+1)(x+5)=0 A. x = 2 or x = 3 B. x = 11 or x = 1 C. x = 1 or x = 5 D. x = -1 or x = -5

OpenStudy (anonymous):

x+1=0 x=-1 x+5=0 x=-5

OpenStudy (anonymous):

D. Because (-1+1)=0 so.. -5 + 5 = 0. So the final answer is D.